### v1

@@ 2 Related work
Optimization-based motion estimation. In the early days of the field, motion estimation results are mostly obtained by optimizing the pairwise non-parametric constraint equations [1, 5, 17], which are further improved by optimizing motion globally over an entire video. Particle video [31] generates a collection of semi-dense long-range trajectories based on initial optical flow fields, while ParticleSfM [45] takes a different approach by optimizing long-range rigid-motion derived from pairwise optical flows within a Structure-from-Motion (SfM) framework. By optimizing a globally consistent canonical volume and coordinate transformation with reliable optical flows, OmniMotion [38] tries to deal with the non-rigid motion of dynamic objects. However, the results obtained by this method are not yet satisfactory.
Matching-based motion estimation. Developments in deep learning have set off a wave of motion estimation by matching, resulting in two dominant approaches: feature matching [39, 19, 36, 24, 21] and dense optical flow [12, 18, 34, 35, 41]. In feature matching, for example, by designing the pretext tasks like frame reconstruction [36, 20], cycle-consistent tracking [39, 19], contrastive learning [42], and masked autoencoders [15, 8], dense representations are learned for matching in a self-supervised manner. On the other hand, optical flow estimators regard it as a regression problem, and employ the cost volumes to find pixel matching. FlowNet [12] first leverages a rendered video dataset to train a regression model, which further encourages the following methods to improve it by building pyramid cost volume[34], iterative inference[35], and multi-frame prediction[32, 16, 46, 3].
@@ 3 Method
This work presents a decoupled representation for test-time optimization, we name it as DecoMotion (Decomposing Motion). The representation explicitly defines two global quasi-3D volumes and transformation functions for dynamic objects and static scenes. As follows, we will first revisit the OmniMotion [38], and then further introduce the proposed decoupled representation. The specific optimization process will also be presented in the next sections.
@@ 3.1 Representation in OmniMotion
OmniMotion represents the scenes in the video with a canonical 3D volume $G$. Following NeRF [28] , OmniMotion defines a coordinate-based network $F_{\theta}$ that maps any 3D point $u\in G$ in that space to the corresponding color value $c$ and density value $\sigma$. The density indicates the quasi-3D scene geometry of the video.
@@ 3.2 Decoupled representation
As analyzed above, dynamic objects and static scenes are very different in terms of both movement and appearance, a unified framework can not represent them well. Therefore, we explicitly decouple the video into two parts, and optimize representations for static scenes and dynamic objects separately. The final representation is obtained by volumetric fusion. The overview of our method is shown in Figure 1.
@@
We also carefully design the static transformation $\mathcal{T}^{\text{st}}$ and the dynamic transformation $\mathcal{T}^{\text{dy}}$, respectively. For the static scenes, the movement caused by the camera’s rigid-motion is mainly considered, so we model the static mapping by optimizing the parameters of the 3D affine transformation between frames, i.e., $\mathcal{T}^{\text{st}}(\cdot)=A_{\theta}(\cdot)$. On the other hand, the motion trajectories of the object are the result of the superposition of the camera and the object’s own motion, which often corresponds to a more complex non-linear transformation. Therefore, we refer to the method of [38], and additionally utilize the Real-NVP[9] to model the dynamic mapping, i.e., $\mathcal{T}^{\text{dy}}(\cdot)=M_{\theta}(A_{\theta}(\cdot);\phi_{i})$, where $\phi_{i}$ stands for the individually optimized latent code for each frame. Due to the ambiguity of motion decomposition, we find that for static transformation, it is also beneficial to appropriately add some non-linear transformation layers, which will be discussed in the subsequent experimental analysis.
@@ 4 Experiment
To validate the proposed method, we evaluate our method on the TAP-Vid benchmark [10], which is designed to evaluate the performance of long-range motion estimation (point tracking) in the real world. This section first describes the experimental setup including dataset, network, training, and evaluation details. After that, we conduct a detailed ablation study to demonstrate how each module works in DecoMotion. Finally, this section will report the performance comparison with the baseline methods on point tracking to further prove the effectiveness of the proposed method for motion estimation.
@@ 4.1 Implementation details
The parameters corresponding to the affine transformation $A_{\theta}$ are predicted by the 2-layer MLP. Meanwhile, as discussed in Figure 5, we also incorporate non-linear invertible transformation layers [9] in the static transformation network and set the number of layers $n_{\text{st}}$ to 3. The $M_{\theta}$ in the dynamic transformation network consists of 6 invertible layers. Prior to this, we have 4 frequency position encoding [28] applied to the input coordinates in each layer. The $F^{\text{dy}}_{\theta}$ and $F^{\text{st}}_{\theta}$ use the GaborNet [13] with 3 layers and 512 channels per layer to encode the color, motion confidence, and density. Considering the complexity of the features, we leverage another GaborNet alone to do the feature rendering. Also, a 2-layer fully-connected layer with 256 channels is included to compute the implicit representation $\phi_{i}$ for each video frame, and the normalized time $t_{i}$ is used as the input to obtain the time embedding.
@@ 4.2 Ablation study
We perform detailed ablation studies on point tracking. Considering that the proposed method is a test-time optimization method, to ensure the executability of the ablation experiments, 10 videos are randomly sampled in TAP-Vid-DAVIS [10], and all quantitative and qualitative results are run at a higher resolution (480p).
@@
We first study how each design in our decoupled representation affects the performance of point tracking. We use the original OmniMotion as the baseline and ensure that the same number of parameters in the transformation network are used for fair comparison. The results are shown in Table 1. The $\mathcal{L}^{\text{st}}$, $\mathcal{L}^{\text{dy}}$ and $\mathcal{L}^{\text{cb}}$ represents the loss functions used in each representation optimization. Solely optimizing representation for static scenes substantially harms the performance, the learned static transformation can not deal with the complex motion patterns of dynamic objects. Whereas the proprietary transformation for dynamic objects improves the tracking performance from 69.9% to 73.3%. With the help of the motion segmentation mask, the dynamic transformation can focus on learning the difficult parts, achieving the globally optimal capability of estimating the object motion. However, the dynamic transformation still fails to capture the camera motion. By combining the static part with $\mathcal{L}^{\text{cb}}$ in volume fusion, the tracking accuracy is further improved to 79.0%. The results fully demonstrate the effectiveness of the divide-and-conquer strategy in motion estimation.
@@
4.2.2 Feature rendering.
We analyze the effect of the proposed feature rendering. The results are shown in Table 2. In (a), we regard the DecoMotion without using any appearance rendering loss as the baseline. With the implicit appearance matching in canonical space, both color and feature rendering enhance the performance of motion estimation. Moreover, the proposed feature rendering further boots up the performance from 76.1% to 79.0%. As shown in the last row of Figure 3, the given query points are misled by the distractors in the background. By using feature rendering loss, the query points are accurately matched to the object in another frame, even facing large deformations, which demonstrates the effectiveness of the representing and rendering features. Nevertheless, the feature rendering on static scenes does not help much. We believe this is because of the relatively poor representation ability of the pre-trained features [6] and the slight inter-frame changes in the background.
@@
Besides, in (b), we use different pre-trained features as supervision, compared with FGVC [23] that more specializes in extracting local features, the VFS [42] and DINO [6], which are better at representing objects, show better performance.
@@
We analyze how each design in the transformation network affects the performance. The results are shown in Figure 5. The $n_{\text{st}},n_{\text{dy}}$ represent the number of non-linear invertible layers [9] used in static and dynamic transformations. For static transformation, purely utilizing affine transformation, i.e., $n_{\text{st}}=0$, gives the worst performance. We find the motion confidence prediction is not always perfect. If a dynamic point is incorrectly predicted as a static point, the affine transformation alone can not deal with it well. Thus, we also add some non-linear layers in the static transformation, which improves the performance. However, including more layers does not bring benefits. Besides, we also increase the $n_{\text{dy}}$ in dynamic transformation, which shows a more significant improvement compared with the method without decoupling.
@@ 4.3 Comparisons on point tracking
We select several representative videos for inference and give the visualization results. DecoMotion produces more accurate and stable pixel correspondences over video frames. For example, in the left case of Figure 6, the dancing girl presents a very complex movement with large deformations, the baseline methods either lose track or get imprecise results, while DecoMotion can consistently track points accurately, which validates the superior ability of DecoMotion to handle the motion of dynamic objects.
@@ 1 More training details
We include more training details here. The base learning rates for the density/color network $F^{\text{st}}_{\theta},F^{\text{dy}}_{\theta}$, the transformation network $M_{\theta}$, and the time embedding MLP are set to $3\times 10^{-4},1\times 10^{-4}$, and $1\times 10^{-3}$, respectively. For loss weight, $\lambda_{1}^{\text{dy}}/\lambda_{2}^{\text{dy}}$ for dynamic objects are set to 1/10, $\lambda_{1}^{\text{st}}/\lambda_{2}^{\text{st}}$ for static scenes are set to 1/10, $\lambda_{1}^{\text{dy}}$ for final representation is set to 1. In the data-preprocessing stage, we use Isomer [44] to obtain the motion segmentation mask. To properly leverage Isomer to process videos of TAP-Vid [10] benchmark for comparisons with state-of-the-art point trackers, we align with the initial training configuration of Isomer [44] by upsampling the original video with a resolution of $256\times 256$ to $384\times 512$. Note we accomplish it by bilinear interpolation to ensure no more information than expected is used in data-preprocessing and training.

### v2

@@ 2 Related work
Optimization-based motion estimation. In the early days, motion estimation results are mostly obtained by optimizing the pairwise non-parametric constraint equations [1, 5, 15], which are improved by optimizing motion globally over an entire video. Particle video [29] generates a collection of semi-dense long-range trajectories based on initial optical flows, while ParticleSfM [43] optimizes long-range rigid-motion derived from pairwise optical flows within a Structure-from-Motion (SfM) framework. By optimizing a globally consistent canonical volume and coordinate transformation with reliable optical flows, OmniMotion [36] tries to deal with the non-rigid motion of dynamic objects. However, the results obtained by this method are not yet satisfactory.
Matching-based motion estimation. By matching between video frames, there are two dominant approaches: feature matching [37, 17, 34, 22, 19] and dense optical flow [11, 16, 32, 33, 39]. In feature matching, for example, by designing the pretext tasks like frame reconstruction [34, 18], cycle-consistent tracking [37, 17], and contrastive learning [40], dense representations are learned for matching in a self-supervised manner. On the other hand, optical flow estimators regard it as a regression problem, and employ the cost volumes to find pixel matching. FlowNet [11] first leverages a rendered video dataset to train a regression model, which further encourages the following methods to improve it by building pyramid cost volume[32], iterative inference[33], and multi-frame prediction[30, 14, 44, 3].
@@ 3 Method
This work presents a decoupled motion representation, we name it as DecoMotion (Decomposing Motion). As follows, we will first revisit the OmniMotion [36], and then further introduce the proposed decoupled representation. The specific optimization process will also be presented in the next sections.
@@ 3.1 Representation in OmniMotion
OmniMotion represents the scenes with a canonical 3D volume $G$. Following NeRF [26] , OmniMotion defines a coordinate-based network $F_{\theta}$ that maps any 3D point $u\in G$ to the corresponding color $c$ and density value $\sigma$. The density indicates the quasi-3D scene geometry of the video.
@@ 3.2 Decoupled representation
As analyzed above, dynamic objects and static scenes are different in terms of movement and appearance. Therefore, we explicitly decouple the video into two parts, and optimize representations for static scenes and dynamic objects separately. The final representation is obtained by volumetric fusion. The overview of our method is shown in Figure 1.
@@
We also carefully design the static transformation $\mathcal{T}^{\text{st}}$ and the dynamic transformation $\mathcal{T}^{\text{dy}}$, respectively. For the static scenes, the movement caused by the camera’s rigid-motion is mainly considered, so we model the static mapping by optimizing the parameters of the 3D affine transformation between frames, i.e., $\mathcal{T}^{\text{st}}(\cdot)=A_{\theta}(\cdot)$. The motion trajectories of the object are the result of the superposition of the camera and the object’s own motion, which often corresponds to a more complex non-linear transformation. Therefore, we refer to the method of [36], and additionally utilize the Real-NVP[8] to model the dynamic mapping, i.e., $\mathcal{T}^{\text{dy}}(\cdot)=M_{\theta}(A_{\theta}(\cdot);\phi_{i})$, where $\phi_{i}$ stands for the individually optimized latent code for each frame. Due to the ambiguity of motion decomposition, we find that for static transformation, it is also beneficial to appropriately add some non-linear transformation layers, which will be discussed in the subsequent experimental analysis.
@@ 4 Experiment
To validate the proposed method, we evaluate our method on the TAP-Vid benchmark [9], which is designed to evaluate the performance of long-range motion estimation (point tracking) in the real world. This section first describes the experimental setup including dataset, network, training, and evaluation details. After that, detailed ablation studies and performance comparisons with the baseline methods will be presented to prove the effectiveness of the proposed method for motion estimation.
@@ 4.1 Implementation details
The parameters corresponding to the affine transformation $A_{\theta}$ are predicted by the 2-layer MLP. Meanwhile, as discussed in Figure 4, we also incorporate non-linear invertible transformation layers [8] in the static transformation network and set the number of layers $n_{\text{st}}$ to 3. The $M_{\theta}$ in the dynamic transformation network consists of 6 invertible layers. The $F^{\text{dy}}_{\theta}$ and $F^{\text{st}}_{\theta}$ use the GaborNet [12] with 3 layers and 512 channels per layer to encode the color, motion confidence, and density. Considering the complexity of the features, we leverage another GaborNet alone to do the feature rendering. Following OmniMotion [36], a 2-layer fully-connected layer with 256 channels is included to compute the implicit representation $\phi_{i}$ for each video frame, and the normalized time $t_{i}$ is used as the input to obtain the time embedding.
@@ 4.2 Ablation study
We perform detailed ablation studies on point tracking. Considering DecoMotion is a test-time optimization method, to ensure the executability of the experiments, a subset of TAP-Vid-DAVIS [9] (480p) is randomly sampled.
@@
We first study how each design in our decoupled representation affects the performance of point tracking. The original OmniMotion is set as the baseline. The results are shown in Table 1. The $\mathcal{L}^{\text{st}}$, $\mathcal{L}^{\text{dy}}$ and $\mathcal{L}^{\text{cb}}$ represents the loss functions used in each representation optimization. Solely optimizing representation for static scenes substantially harms the performance, the learned static transformation can not deal with the complex motion patterns of dynamic objects. Whereas the proprietary transformation for dynamic objects improves the tracking performance from 69.9% to 73.3%. With the help of the motion segmentation mask, the dynamic transformation can focus on learning the difficult parts, achieving the globally optimal capability of estimating the object’s motion. However, the dynamic transformation still fails to capture the camera motion. By combining the static part with $\mathcal{L}^{\text{cb}}$ in volume fusion, the tracking accuracy is further improved to 79.0%. The results demonstrate the effectiveness of the divide-and-conquer strategy in motion estimation.
@@
4.2.2 Feature rendering.
We analyze the effect of the proposed feature rendering. The results are shown in Table 2. In (a), we regard the DecoMotion without using any appearance rendering loss as the baseline. With the implicit appearance matching in canonical space, both color and feature rendering enhance the performance of motion estimation. Moreover, the proposed feature rendering further boots up the performance from 76.1% to 79.0%. As shown in the last row of Figure 2, the given query points are misled by the distractors in the background. By using feature rendering loss, the query points are accurately matched to the object in another frame, even facing large deformations. Nevertheless, the feature rendering on static scenes does not help much. We believe this is because of the relatively poor representation ability of the pre-trained features [6] and the slight inter-frame changes in the background. Besides, in (b), we use different pre-trained features as supervision, compared with FGVC [21] that more specializes in extracting local features, the VFS [40] and DINO [6], which are better at representing objects, show better performance.
@@
In Figure 4, the $n_{\text{st}},n_{\text{dy}}$ represent the number of non-linear invertible layers [8] used in static and dynamic transformations. For static transformation, purely utilizing affine transformation, i.e., $n_{\text{st}}=0$, gives the worst performance. We find the motion confidence prediction is not always perfect. If a dynamic point is incorrectly predicted as a static point, the affine transformation alone can not deal with it well. Adding some non-linear layers in the static transformation improves the performance. However, including more layers does not bring benefits. Besides, increasing the $n_{\text{dy}}$ in dynamic transformation shows a more significant improvement compared with the method without decoupling.
@@ 4.3 Comparisons on point tracking
We select several representative videos for inference. DecoMotion produces more accurate and stable pixel correspondences over frames. For example, in the left case of Figure 5, the dancing girl presents a very complex movement with large deformations, the baseline methods either lose track or get imprecise results, while DecoMotion can consistently track points accurately, which validates the superior ability of DecoMotion to handle the motion of dynamic objects.
@@ 5 Conclusion
The supplementary material contains: 1) more details about the training; 2) the demo video of feature rendering; 3) the demo video of appearance decomposition; 4) the demo video consists of more qualitative results of point tracking. 5) more results of tracking on multiple dynamic objects. 6) the limitations of DecoMotion.
@@ 1 More training details
We include more training details here. The base learning rates for the density/color network $F^{\text{st}}_{\theta},F^{\text{dy}}_{\theta}$, the transformation network $M_{\theta}$, and the time embedding MLP are set to $3\times 10^{-4},1\times 10^{-4}$, and $1\times 10^{-3}$, respectively. For loss weight, $\lambda_{1}^{\text{dy}}/\lambda_{2}^{\text{dy}}$ for dynamic objects are set to 1/10, $\lambda_{1}^{\text{st}}/\lambda_{2}^{\text{st}}$ for static scenes are set to 1/1, $\lambda_{1}^{\text{cb}}$ for final representation is set to 1. In the data-preprocessing stage, we use Isomer [42] to obtain the motion segmentation mask. To properly leverage Isomer to process videos of TAP-Vid [9] benchmark for comparisons with state-of-the-art point trackers, we align with the initial training configuration of Isomer [42] by upsampling the original video with a resolution of $256\times 256$ to $384\times 512$. Note we accomplish it by bilinear interpolation to ensure no more information than expected is used in data-preprocessing and training.